### v1

# Key Words.: Galaxies: clusters: general – Galaxies: clusters: individual: Abell 426 – X-rays: galaxies: clusters – Radio continuum: galaxies – Galaxies: clusters: intracluster medium – Galaxies: kinematics and dynamics1 Introduction

Perseus cluster in its X-ray entirety with SRG/eROSITA et al. I. Merger and Radio-Uroboroses.
The Perseus cluster (Abell 426) is a nearby massive galaxy cluster that spans several degrees. We combined SRG/eROSITA, XMM-Newton, and Chandra data to get a complete coverage of this cluster in X-rays up to and beyond, although at the largest radii, spatial non-uniformities of X-ray background and foreground dominate. While the Perseus central part represents a canonical cool-core structure with clear signs of AGN Feedback, the outskirts, in turn, serve as a convincing example of a merger-perturbed system. X-ray data suggest that IC310 is the main galaxy of a subcluster that merges with Perseus over the past . Overall, this configuration resembles the merger between the Coma cluster and the NGC4839 group. It is statistically more likely to find a merging group near the apocenter of its orbit. Therefore, it is not surprising that IC310 in Perseus has a relatively small velocity relative to the main cluster, similarly to NGC4839 in Coma.
Perseus also hosts a high-velocity radio galaxy, NGC1265 (line-of-sight velocity is almost twice the virial velocity of the main cluster), which is known for its spectacular radio tail. Unless this galaxy has been accelerated by a time-variable potential associated with the merger, it has to move almost along the line of sight through the entire cluster, which would be a rare, but not a truly exceptional configuration. Both galaxies, IC310 and NGC1265, have remarkable radio tails with sharp bends that are reminiscent of a ”snake biting its tail”. We speculate that these curious shapes are natural consequences of their (different) orbits in Perseus. For IC310, the proximity to the apocenter and the reversal of its radial velocity might play a role. For NGC1265, the nearly line-of-sight motion coupled with the gas motions in the merging system might be important.
Key Words.:
Galaxies: clusters: general – Galaxies: clusters: individual: Abell 426 – X-rays: galaxies: clusters – Radio continuum: galaxies – Galaxies: clusters: intracluster medium – Galaxies: kinematics and dynamics1 Introduction
The Perseus cluster (Abell 426) is the X-ray brightest and one of the best-studied clusters. Its core properties were used to develop the concept of ”AGN feedback” (e.g., Boehringer et al., 1993; Churazov et al., 2000; Fabian et al., 2003; Zhuravleva et al., 2015). In the core and at larger scales, X-ray images show signs of perturbations, including a sequence of sloshing patterns located at different distances from the center (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Zhang et al., 2020; Sanders et al., 2020; Walker et al., 2022; Zhu et al., 2021), (see, also Markevitch & Vikhlinin, 2007; Zuhone & Roediger, 2016, for reviews on merger-induced perturbations).
The complications associated with the X-ray analysis of the Perseus cluster are caused by a) proximity to the Galactic Plane (hence, significant photoelectric absorption) and the large angular size, which complicates mapping the entire cluster with telescopes having a relatively small field of view. The latter problem can be overcome with all-sky surveys. In this study, we report SRG/eROSITA (Sunyaev et al., 2021; Predehl et al., 2021) observations of the Perseus cluster up to its virial radius. We combine eROSITA data with the publicly available Chandra and XMM-Newton data to achieve high angular resolution in the core of the cluster and a uniform coverage in the outskirts.
In this study, we adopted the following basic properties of the Perseus cluster: (Hitomi Collaboration et al., 2016). For , this redshift translates to the angular diameter distance of 74.1 Mpc; . We adopt111See also, Meusinger et al. (2020) for another set of parameters which imply a larger mass of the Perseus cluster. and (Simionescu et al., 2011) and the corresponding circular velocity . We also assume that can be used in place of the virial radius (a parameter that enters the mass model).
2 X-ray data
For X-ray analysis, we use the data accumulated by the eROSITA (Predehl et al., 2021) telescope onboard the SRG X-ray observatory (Sunyaev et al., 2021) in the course of its all-sky survey. Initial reduction and processing of the data were performed using standard routines of the eSASS software (Brunner et al., 2018; Predehl et al., 2021), while the imaging and spectral analysis were carried out with the background models, vignetting, point spread function (PSF) and spectral response function calibrations built upon the standard ones via slight modifications motivated by results of calibration and performance verification observations (e.g. Churazov et al., 2021; Khabibullin et al., 2023).
For XMM-Newton and Chandra, we used a subset of observations accumulated over the lifetime of these missions. The lists of observations are given in the Appendix A.
2.1 Large-scale image
A 10x10 degrees patch of the sky centered at the Perseus cluster is shown in Fig. 1. The eROSITA 0.4-2.3 keV image illustrates the connection of Perseus to the large-scale structure, in particular, two bright clusters 2A0251+413 () and CIZAJ0300.7+4427 (), which trace filamentary structures in the vicinity of the main cluster. The same structures are visible in the PLANCK -parameter map (Fig. 1, middle panel), where is proportional to the line-of-sight integrated electron pressure and can be derived from distortions of the Cosmic Microwave Background (see Zeldovich & Sunyaev, 1969, where the formulae describing these distortions were found).
The location of the Perseus cluster close to the Galactic Plane introduces some biases in the X-ray and -maps. The extinction derived from the Bayestar data (Green et al., 2019) varies from 0.1 to 3.6 across the image. As a result, significant variations of photoelectric absorption modulate the X-ray image, and the dust-correlated structures are also visible in the -map (see, e.g., the bottom-left corner of the images). Assuming that can be converted to equivalent hydrogen column density as (e.g., Predehl & Schmitt, 1995; Güver & Özel, 2009) and considering the contribution of atomic gas (traced by HI, ), we can expect a modulation of the Perseus X-ray flux in the 0.4-2.3 keV up to within a circle of 1.4 degrees. This estimate is done for the APEC spectrum with a temperature of 5 keV. The total attenuation factor of the cluster flux (relative to ) is about 0.55.
2.2 2MASS
It is useful to compare the X-ray images with the distribution of galaxies with redshifts comparable to that of the cluster. To this end, we used the 2MASS Redshift Survey (Huchra et al., 2012) and selected galaxies with line-of-sight recession velocity in the range from to . The distribution of these galaxies is shown in Fig. 2. The left (20x20 degrees) image shows clearly the famous large-scale filaments in the vicinity of Perseus, which are part of the Perseus-Pisces supercluster. The most prominent two filaments, relevant for this study, are marked with black lines. One of them is nearly horizontal in equatorial coordinates and is dominated by galaxies with redshifts similar to the main cluster. Another one is diagonal in this plot, and its upper end is dominated by galaxies with larger recession velocities (blue colors). The right image (2x2 degrees) is the zoomed version of the same plot. The chain of galaxies between NGC1275 and IC310 appears to align with the former filament, while NGC1265 might be part of the latter.
2.3 X-ray composite image
Focusing on the X-ray emission inside the Perseus virial radius, we would like to complement the eROSITA coverage of the entire field with higher angular resolution of Chandra and XMM-Newton in the Perseus central region. We first generated separate background-subtracted, exposure and vignetting-corrected X-ray images for eROSITA (0.4-2.3 keV; 4” pixels), XMM-Newton (0.5-3.5 keV; 2” pixels), and Chandra (0.5-3.5 keV, 1” pixels). These images have been combined in two steps. In the first step, the mean surface brightness in a circle centered at NGC1275 has been calculated, and the normalization of images was adjusted to compensate for slight differences in the energy bands and effective areas. The resulting images were rebinned to 2” resolution and co-added using a position-dependent weighting factor. This factor favors the Chandra images in the core and the eROSITA image at the largest distances from NGC1275. The resulting 2”-resolution, degrees image is shown in Fig. 3. The inset shows the central 1% of the same image, where Chandra’s resolution captures familiar structures of the X-ray emission in the Perseus core.
2.4 Departures from a symmetric model
The known elongation of the Perseus X-ray emission along the East-West direction is already visible in Fig. 3. To emphasize further the departures from spherical symmetry, we fitted the simplest beta model to the radial X-ray surface brightness profile. The brightest compact sources in the field were excised prior to fitting the model. The profile and the best-fitting beta model are shown in the Appendix B. The dynamic range covered by the profile from the inner to approaches a factor of . Overall, the beta model gives a reasonable (for our purposes) approximation of the Perseus X-ray surface brightness. We note in passing that at radii in the range 1-3 degrees, a stray light can give a non-negligible or even dominant contribution to the surface brightness profile. However, to reveal asymmetric features in the outskirts of the Perseus cluster, this is not required as long as the flux from the cluster core dominates the stray light and, therefore, the model remains symmetric, even though the amplitude of deviations can be affected.
Fig. 4 shows the X-ray image divided by the model. The model is the sum of the cluster emission (the -model described above) and the constant sky background. Once again, the core shows the familiar pattern of X-ray cavities and several spiral-like structures, while on larger scales, the dominant features are two large regions of excess emission on both sides of the core, which are roughly aligned with the East-West direction. Many of these structures (or parts of them) have already been seen in previous X-ray observations (e.g., Branduardi-Raymont et al., 1981; Schwarz et al., 1992; Ettori et al., 1998; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Tamura et al., 2014; Walker et al., 2022; Zhu et al., 2021), but the complete and uniform coverage of eROSITA reveals this pattern in its entirety.
3 Discussion
3.1 Merger geometry and timing
The dynamic state of the Perseus cluster (outside the very core) is a subject of many studies over several decades, (see, e.g., Branduardi-Raymont et al., 1981; Kent & Sargent, 1983; Andreon, 1994; Churazov et al., 2003; Fabian et al., 2011; Simionescu et al., 2012; Bellomi et al., 2024; Sanders et al., 2020; Kang et al., 2024; HyeongHan et al., 2025, among others) with an overall conclusion that the cluster is neither fully relaxed nor in a state of a major merger.
The residual X-ray images (Figs. 4 and 6) show that the excess X-ray emission is well centered at IC310. IC310 is an S0 galaxy that hosts an AGN (BL Lac) that is visible from the radio to TeV bands (e.g. Graham et al., 2019). Due to the X-ray bright AGN, the constraints on the diffuse X-ray emission in the vicinity of IC310 are marginal, although the presence of the excess emission (or substructure) on somewhat larger scales has been found (see Schwarz et al., 1992; Sato et al., 2005; Dunn et al., 2010; Simionescu et al., 2012). eROSITA data support the interpretation that IC310 is the center of a subcluster merging with the Perseus cluster. Further support for this statement comes from a chain of bright galaxies that apparently ”connect” NGC1275 and IC310 (see Fig. 2). Yet another piece of evidence comes from radio data, which we discuss in Sect. 3.3.
IC310 is located (in projection) to the West of NGC1275. This galaxy has a relatively low line-of-sight velocity compared to the cluster , which is about half of the circular speed in the main cluster. In the merger scenario, the perturbing subcluster spends, on average, more time near the apocenter than near the pericenter. If this is the case, the low velocity is expected, and no strong constraints on the true 3D position of IC310 can be placed222 A similar merger configuration is suggested for the Coma cluster, where the NGC4839 group (merging subcluster) has a small line-of-sight velocity relative to the main cluster (e.g. Lyskova et al., 2019).. At the same time, the lack of obvious gradients in redshifts of galaxies along the main filament (Fig. 2) suggests that the merger direction is not far from the sky plane, in which case the 3D velocity might be higher. While IC310 itself does not directly constrain the ”age” of the merger, one can hope to use a substructure in the ICM distribution that retains some memory of past perturbations. Those include merger shocks and sloshing of the core gas.
The association of the prominent asymmetric perturbations (beyond the region strongly dominated by AGN Feedback) in the Perseus cluster (see, e.g., Schwarz et al., 1992; Simionescu et al., 2012) with gas sloshing and the Brunt-Vaisala frequency became clear (see Fig. 12 in Churazov et al., 2003) once the first reliable ICM temperature maps with sufficient angular resolution became available. These maps helped to differentiate between shocks, AGN-inflated bubbles, and contact discontinuities. Using numerical simulations, Roediger et al. (2024) calibrated the effective ”propagation velocity” of the contact discontinuity pattern for a Perseus-like cluster. This relation predicts that , where is the ICM sound speed and is the time since the initial perturbation (see Roediger et al., 2024, for discussion of limitations of this expression). Assuming that this relation applies to the prominent excess to the East of the center (see Fig. 4), we can estimate the age of this structure. Adopting and the distance from the Perseus core (= NGC1275) of 700 kpc, one gets 333This is somewhat shorter than suggested by HyeongHan et al. (2025), where the weak-lensing-based mass concentration associated with NGC1264 (a galaxy NGC1262 some 400 kpc to the West of NGC1275) was suggested as a center of the local mass concentration..
A weak shock propagating with velocity during this time would be at the distance of , well beyond the region studied here. The characteristic time for the subcluster motion (= IC310 in our model) is probably a multiple of the viral time . Therefore, it is plausible that the merger shock associated with the pericenter passage that initiated the most prominent contact discontinuity to the East of NGC1275 is long gone, while the IC310 group could be currently at its second (or third) apocenter (see, also, simulations of Bellomi et al., 2024). Further support for the ”near apocenter” scenario is discussed in Sect. 3.3.
The merger scenario with the excess mass westwards of NGC1275 is also supported by the recent weak lensing data (HyeongHan et al., 2025), albeit the suggested centroid (near the galaxy NGC1264) is a factor of closer to the cluster center than IC310. The most recent Euclid data (Kluge et al., 2025; Marleau et al., 2025) also show the excess of the intracluster light and dwarf galaxy to the west of the core.
3.2 Substructure near the virial radius
At larger scales, comparable to the virial radius, some structures have been reported based on the X-ray data of XMM-Newton and Suzaku. In particular, (Walker et al., 2022) suggested two X-ray surface brightness edges at 1.2 and 1.7 Mpc to the west of the core, which were interpreted as cold fronts. Zhang et al. (2020) argued that these cold fronts (= contact discontinuities) at very large distances from the center might not be the sloshing structures but rather the result of the collision between the accretion shock and a ’runaway’ merger shock (see also Birnboim et al., 2010). The analysis of the additional Suzaku data also revealed a discontinuity near (to the North-West from the core), but from the projected temperature profile, it was concluded that this structure is an shock (Zhu et al., 2021). The eROSITA broad-band data confirm the presence of the substructure in this direction, although the analysis of the spectral characteristics will be reported elsewhere.
The smoothed eROSITA+Chandra+XMM-Newton residual map444Only the eROSITA data are smoothed in this combined image to retain the good angular resolution in the core. is shown in Fig. 6. This image does show the excess emission westwards to the core extending to the virial radius. Some possible structures near are also seen in the NW direction. We defer the interpretation of these features for future studies. Here, we only argue that in the minimalists’ scenario, these features are plausibly associated with the same merger event. We also note that there is marginal evidence for other large-scale features in the residual X-ray image (Fig. 4), e.g., an arc-like feature running almost vertically through NGC1265. However, more data are needed to assess its reality.
3.3 Two prominent radio galaxies
Apart from the central Perseus galaxy, famous NGC1275, the Perseus cluster hosts many prominent radio galaxies (see, e.g., Gendron-Marsolais et al., 2020; van Weeren et al., 2024, for recent VLA nd LOFAR maps). Here, we focus on two of them - NGC1265 and IC310. The LOFAR radio image (a combination of and resolution maps from van Weeren et al. 2024) superposed on the X-ray excess emission image is shown in Fig. 7.
3.3.1 IC310
As mentioned above, IC310 appears to be at the center of the subcluster/group that merges with the main cluster. In terms of IC310 offset position ( from the core) and low line-of-sight velocity ( relative to NGC1275)555Quoted relative line-of-sight velocities are calculated as ., this galaxy is reminiscent of the NGC4839 group in the Coma cluster. The low velocity might be related to the proximity of the group to the apocenter, hence low 3D and line-of-sight velocities. While IC310 is originally classified as a narrow-tail galaxy, which would be more consistent with a high relative velocity of the galaxy relative to ICM (Begelman et al., 1979), it might be a special case of a radio galaxy seen nearly edge-on Sijbring & de Bruyn (1998); Gendron-Marsolais et al. (2020). This interpretation is supported by the blazar-like properties of the nucleus (e.g. Kadler et al., 2012; Aleksić et al., 2014), with a pc-scale jet making a relatively small angle to the line of sight666Another interesting detail discussed by Kadler et al. (2012) is that the orientation of the pc-scale jet is similar to that of the kpc-long tail.. The most recent LOFAR data revealed a much more extended radio tail of IC310 (van Weeren et al., 2024). The tail makes a sharp turn towards the Perseus center (see Fig. 7). This morphology could be explained by a merger scenario, where IC310 has already gone through the pericenter and the apocenter and is now moving back towards the core. In this case, the ”reversal” in the far tail direction is explained by the so-called slingshot effect (Lyskova et al., 2019; Sheardown et al., 2019; Zhang et al., 2019), when the ICM shifted backwards by the ram pressure, swings to the opposite side when the motion of the halo slows down, and then reverses near the apocenter. This configuration leads to a loop-like geometry of the tail that we dubbed ”radio-Uroboros”.
Overall, the optical, radio, and X-ray data support the IC310-related merger scenario, happening along the direction of the major East-West filament crossing the Perseus cluster.
3.3.2 NGC1265
Unlike IC310, NGC1265 has a very high velocity relative to NGC1275 () and it lacks obvious large-scale X-ray emission around the X-ray bright core. While there are even higher velocity galaxies in the same region (plausibly associated with another large-scale filament, see Fig. 2), the radio tail suggests interaction with the Perseus ICM. The tail identified in Ryle & Windram (1968); Miley et al. (1975), led to the classification of NGC1265 as a Narrow-angle-tail galaxy, although at higher resolution, the core rather looks like wide-angle-tail galaxy (see a discussion of tags and boxes in Rudnick, 2021). The extent of its long and bent tail (projected length ) was revealed with WSRT observations (Sijbring & de Bruyn, 1998) and more recently imaged in great detail (Gendron-Marsolais et al., 2020; van Weeren et al., 2024).
Remarkably, despite being very different from IC310, the radio tail of NGC1265 also qualifies for the ”radio-Uroboros” tag (see Fig. 7). The statistic of radio tails’ direction suggests that for a bare galaxy, especially moving with a very high velocity, the gas feeding the central AGN can be stripped during the first infall (e.g., Roberts & Parker, 2020; Lal et al., 2022), NGC1265 appears to be a newcomer to the Perseus cluster. Yet, it has a very extended tail (probably a factor of a few longer than its apparent projected size). In Pfrommer & Jones (2011), the curious properties of the NGC1265 tail are explained as due to the galaxy crossing an accretion shock near the virial radius, with the galaxy located at the near side of the cluster. While this is a plausible scenario, we note that the ICM in the Perseus can be significantly perturbed by the ongoing merger, and for a galaxy on a nearly straight trajectory, dictated by its high velocity, the ICM motions can perturb the radio tail left by NGC1265. We, therefore, consider below a simple model with NGC1265 at the far side (along the line of sight) of the main cluster. This model broadly corresponds to scenarios c) and d) in Sijbring & de Bruyn (1998).
3.3.3 Test particle (NGC1265) in a static potential
As the starting point, we consider a test particle moving in a static potential well having an NFW mass distribution (Navarro et al., 1997). We further assume a zero total energy for this particle, i.e, . We chose a coordinate system with and in the sky plane, and along the line-of-sight. We aim to reproduce the line-of-sight velocity and the projected distance of NGC1265. The requirement to have zero total energy implies that the 3D distance of the particle from the cluster center is less than , while the lower limit is set by the observed projected distance from the cluster center, i.e., (see Fig. 8). If the marginally bound particle is close to , it has to move exactly along the line of sight. In the opposite case, the angle to the line of sight is set by the relation between the local 3D velocity (set by the 3D radius) and the . For the limiting case of , this angle is about degrees. For any realistic 3D position (further away than ), this angle will be smaller. If the overall shift of the tail eastwards of the current position of NGC1265 reflects the large-scale trajectory of the galaxy, it should be crossing the main cluster on a slightly curved trajectory. Given the high velocity of NGC1265, the curved trajectory requires the galaxy to cross the core of the cluster. A sample of such trajectories is shown in Fig. 9. They all correspond to the same (current) projected distance, line-of-sight velocity, and the velocity direction in the sky plane and differ in the magnitude of the 3D velocity. This is, of course, a small subsample of possible trajectories (given direct observational constraints and our assumption that the galaxy is currently on the far side of the cluster). In this model, we assume that gas motions in the region between NGC1275 and IC310 further disturb the tail trajectory to give it the observed shape. No attempt was made to get constraints on the detailed properties of the tail, given that the density and pressure of the ICM change significantly along the galaxy trajectory, and the power of the NGC1265 jets can vary too.
The above model does not consider constraints coming from the uncertainties associated with the Faraday Rotation measurement in the direction of NGC1265 (see, e.g., de Bruyn & Brentjens, 2005; Brentjens, 2011), that might predominantly come from the Milky Way rather than from the cluster (see Pfrommer & Jones, 2011, for a scenario with NGC1265 on the near side of the cluster).
Another limitation comes from the assumption of a static potential of the main cluster. In reality, peculiar initial velocities and time variations of the potential can play a role. For a merging cluster such as Perseus, the latter effect might be important (e.g., Sales et al., 2007; Carucci et al., 2014). In this case, the trajectory of the galaxy might be more complicated. We do not discuss this scenario further.
Overall, we see several attractive features of the model described above, but we can not exclude other scenarios.
3.3.4 Line-of-sight velocities in simulations
While the NGC1265 trajectories discussed in Sect. 3.3.3 were derived assuming zero binding energy of the galaxy in a static NFW potential, one can pose a question about how often we expect to find a galaxy given its projected distance and the line-of-sight velocity (with both quantities expressed in terms of virial properties of the main halo). Since the answer depends on the mass accretion rate of the main halo and on the importance of other effects such as dynamic friction (for massive subhalos), it can be best addressed by looking at the results of numerical simulations. Here, we use a publicly available catalog of halos in Magneticum777www.magneticum.org simulations set (Hirschmann et al., 2014; Dolag et al., 2016, 2025), see also Aung et al. 2023.
These simulations were performed with the TreePM/SPH code GADGET-3 Springel (2005); Beck et al. (2016); Groth et al. (2023). They take into account many complex non-gravitational physical processes (cooling, merging of galaxies and galaxy clusters, shock waves, and detailed galaxy formation physics, including the treatment of black holes) which determine the evolution of large-scale systems and affect their observational properties. More details of the physics provided in the simulation can be found in Section 2 of Biffi et al. (2022) and references therein. We take advantage of the haloes catalogue extracted from Box2/hr simulation box at its terminal redshift. Its comoving volume is equal to ( is the Hubble constant), containing mass resolution elements (particles). The masses of dark matter and gas particles are equal to and , respectively, and the following cosmological parameters are adopted: the total matter density ( baryons), the cosmological constant , the Hubble constant km/s/Mpc (i.e. ), the index of the primordial power spectrum , the overall normalization of the power spectrum (Komatsu et al., 2011).
For each halo more massive than , we identified all subhalos. The distances (from the halo to subhalos) and velocities have been rescaled to and of the main halo888We used the ”500c”-based quantities (rather than ”200c”) since they are provided in the catalogs. is close to anyway, while the relation between and changes from 0.63 to 0.66 for concentration parameter between and . This information was used to construct the probability density function of the line-of-sight velocity as a function of projected distance. Namely, each subhalo, located at a distance of from the main halo, was accounted for when considering all possible viewing angles corresponding to projected distances from to and the corresponding solid angles. We also set the sign of the line-of-sight velocity based on the sign of the radial component of the velocity. Negative values correspond to subhalos currently moving towards the main halo999Such information is not available in observations.. The resulting PDF is shown in Fig. 10 for three values of projected distances. As expected, at large projected distances, the PDF is asymmetric due to the ”first-infall” region of the phase space, while at smaller distances this difference becomes less pronounced. The two vertical dashed lines show the line-of-sight velocity of NGC1265, which is . The projected distance of NGC1265 from the Perseus core is (the black curve in Fig. 10. The probability of finding a galaxy at this projected distance with the line-of-sight velocity larger than is less than 1% (actually, close to 0.1%). This makes NGC1265 a very rare object, but not truly exceptional.
We apply the data from the same catalogues to the IC310 case, but this time we fix the projected distance and plot in Fig.11 the velocity PDF separately for low mass galaxies and more massive objects (groups), implicitly assuming that IC310 is indeed the central galaxy of a subcluster and belongs to the ”massive” subsample. This separation was based on the stellar mass, because this quantity is a somewhat better-defined (and conserved) quantity than the total mass. The high mass bin effectively contains all massive halos, i.e., clusters themselves, but they do not contribute to the plotted distribution since they have zero offset from the parent cluster. The projected distance of IC310 is , while its velocity is (see two vertical green lines in Fig.11). For the low mass sample (the black curve), the distribution is symmetric, and IC310 (if it is an individual galaxy rather than the center of the group) is in the central part of the distribution and has an equal probability of being infalling and outgoing. For the high mass case (the red line), the observed line-of-sight velocity is also consistent with both cases, while a factor of larger observed velocity would already be in tension with the outgoing scenario. Based on these estimates, we concluded that the infalling and outgoing scenarios can not be excluded for IC310, even if it is a group of galaxies.
4 Conclusions
We combined Chandra, XMM-Newton, and eROSITA data to make a complete image of the Perseus cluster from the very core to the virial radius. This image shows all the spectacular features that a bona fide cluster might have: clear signs of AGN Feedback in the core, Cold Fronts (contact discontinuities) at intermediate scales, and signatures of a merger at the largest scales.
eROSITA X-ray data suggest that the IC310 galaxy ( west of NGC1275) is at the center of the subcluster/group that perturbed Perseus some 4 Gyr ago. A clear excess of X-ray emission is seen in a Mpc-size region around IC310, implying that the subcluster was able to retain some of its gaseous atmosphere.
IC310 has a rather small recession velocity relative to the cluster. We tentatively attribute this small velocity to the selection effects - it is more likely to find the infalling object near the apocenter than near the pericenter. This is, of course, a qualitative argument. It is supported by the shape of the IC310 radio tail, which bends sharply in the vicinity of the galaxy. Such a sharp turn could arise when the sign of the radial velocity changes near the apocenter. Alternatively, the low velocity of IC310 could reflect the merger direction in the sky plane, given that only a weak gradient in recession velocities is seen in the major large-scale filament running through Perseus.
At larger scales, the excess X-ray emission can be traced westwards up to the virial radius, with some tentative arc-like features in the NW direction. It is plausible that these structures are associated with the same merger.
Finally, we discussed briefly the case of the high-velocity radio galaxy NGC1265, known for its curious radio tail, which also shows sharp bends. In our ”minimalist” model, this galaxy is moving nearly along the line of sight with the total 3D velocity and is currently on the far side of the main cluster. The gas motions induced by IC310 merger might play a role in shaping its tail, although other explanations are possible, too.
Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI), in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. IK and KD acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pr83li (Magneticum).References
- Aleksić et al. (2014) Aleksić, J., Ansoldi, S., Antonelli, L. A., et al. 2014, Science, 346, 1080
- Andreon (1994) Andreon, S. 1994, A&A, 284, 801
- Aung et al. (2023) Aung, H., Nagai, D., Rozo, E., Wolfe, B., & Adhikari, S. 2023, MNRAS, 521, 3981
- Beck et al. (2016) Beck, A. M., Murante, G., Arth, A., et al. 2016, MNRAS, 455, 2110
- Begelman et al. (1979) Begelman, M. C., Rees, M. J., & Blandford, R. D. 1979, Nature, 279, 770
- Bellomi et al. (2024) Bellomi, E., ZuHone, J. A., Weinberger, R., et al. 2024, ApJ, 974, 234
- Biffi et al. (2022) Biffi, V., Dolag, K., Reiprich, T. H., et al. 2022, A&A, 661, A17
- Birnboim et al. (2010) Birnboim, Y., Keshet, U., & Hernquist, L. 2010, MNRAS, 408, 199
- Boehringer et al. (1993) Boehringer, H., Voges, W., Fabian, A. C., Edge, A. C., & Neumann, D. M. 1993, MNRAS, 264, L25
- Branduardi-Raymont et al. (1981) Branduardi-Raymont, G., Fabricant, D., Feigelson, E., et al. 1981, ApJ, 248, 55
- Brentjens (2011) Brentjens, M. A. 2011, A&A, 526, A9
- Brunner et al. (2018) Brunner, H., Boller, T., Coutinho, D., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10699, Space Telescopes and Instrumentation 2018: Ultraviolet to Gamma Ray, ed. J.-W. A. den Herder, S. Nikzad, & K. Nakazawa, 106995G
- Carucci et al. (2014) Carucci, I. P., Sparre, M., Hansen, S. H., & Joyce, M. 2014, J. Cosmology Astropart. Phys, 2014, 057
- Chandran et al. (2023) Chandran, J., Remazeilles, M., & Barreiro, R. B. 2023, MNRAS, 526, 5682
- Churazov et al. (2000) Churazov, E., Forman, W., Jones, C., & Böhringer, H. 2000, A&A, 356, 788
- Churazov et al. (2003) Churazov, E., Forman, W., Jones, C., & Böhringer, H. 2003, ApJ, 590, 225
- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
- Churazov et al. (2021) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021, A&A, 651, A41
- de Bruyn & Brentjens (2005) de Bruyn, A. G. & Brentjens, M. A. 2005, A&A, 441, 931
- Dolag et al. (2016) Dolag, K., Komatsu, E., & Sunyaev, R. 2016, MNRAS, 463, 1797
- Dolag et al. (2025) Dolag, K., Remus, R.-S., Valenzuela, L. M., et al. 2025, arXiv e-prints, arXiv:2504.01061
- Dunn et al. (2010) Dunn, R. J. H., Allen, S. W., Taylor, G. B., et al. 2010, MNRAS, 404, 180
- Ettori et al. (1998) Ettori, S., Fabian, A. C., & White, D. A. 1998, MNRAS, 300, 837
- Fabian et al. (2011) Fabian, A. C., Sanders, J. S., Allen, S. W., et al. 2011, MNRAS, 418, 2154
- Fabian et al. (2003) Fabian, A. C., Sanders, J. S., Allen, S. W., et al. 2003, MNRAS, 344, L43
- Gendron-Marsolais et al. (2020) Gendron-Marsolais, M., Hlavacek-Larrondo, J., van Weeren, R. J., et al. 2020, MNRAS, 499, 5791
- Graham et al. (2019) Graham, J. A., Brown, A. M., & Chadwick, P. M. 2019, MNRAS, 485, 3277
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
- Groth et al. (2023) Groth, F., Steinwandel, U. P., Valentini, M., & Dolag, K. 2023, MNRAS, 526, 616
- Güver & Özel (2009) Güver, T. & Özel, F. 2009, MNRAS, 400, 2050
- Hirschmann et al. (2014) Hirschmann, M., Dolag, K., Saro, A., et al. 2014, MNRAS, 442, 2304
- Hitomi Collaboration et al. (2016) Hitomi Collaboration, Aharonian, F., Akamatsu, H., et al. 2016, Nature, 535, 117
- Huchra et al. (2012) Huchra, J. P., Macri, L. M., Masters, K. L., et al. 2012, ApJS, 199, 26
- HyeongHan et al. (2025) HyeongHan, K., Jee, M. J., Lee, W., et al. 2025, Nature Astronomy, 9, 925
- Kadler et al. (2012) Kadler, M., Eisenacher, D., Ros, E., et al. 2012, A&A, 538, L1
- Kang et al. (2024) Kang, W., Hwang, H. S., Song, H., et al. 2024, ApJS, 272, 22
- Kent & Sargent (1983) Kent, S. M. & Sargent, W. L. W. 1983, AJ, 88, 697
- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
- Kluge et al. (2025) Kluge, M., Hatch, N. A., Montes, M., et al. 2025, A&A, 697, A13
- Komatsu et al. (2011) Komatsu, E., Smith, K. M., Dunkley, J., et al. 2011, ApJS, 192, 18
- Lal et al. (2022) Lal, D. V., Lyskova, N., Zhang, C., et al. 2022, ApJ, 934, 170
- Lyskova et al. (2019) Lyskova, N., Churazov, E., Zhang, C., et al. 2019, MNRAS, 485, 2922
- Markevitch & Vikhlinin (2007) Markevitch, M. & Vikhlinin, A. 2007, Phys. Rep., 443, 1
- Marleau et al. (2025) Marleau, F. R., Cuillandre, J. C., Cantiello, M., et al. 2025, A&A, 697, A12
- Meusinger et al. (2020) Meusinger, H., Rudolf, C., Stecklum, B., et al. 2020, A&A, 640, A30
- Miley et al. (1975) Miley, G. K., Wellington, K. J., & van der Laan, H. 1975, A&A, 38, 381
- Navarro et al. (1997) Navarro, J. F., Frenk, C. S., & White, S. D. M. 1997, ApJ, 490, 493
- Pfrommer & Jones (2011) Pfrommer, C. & Jones, T. W. 2011, ApJ, 730, 22
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Predehl & Schmitt (1995) Predehl, P. & Schmitt, J. H. M. M. 1995, A&A, 293, 889
- Roberts & Parker (2020) Roberts, I. D. & Parker, L. C. 2020, MNRAS, 495, 554
- Roediger et al. (2024) Roediger, E., Vaezzadeh, I., & Nulsen, P. 2024, MNRAS, 529, 563
- Rudnick (2021) Rudnick, L. 2021, Galaxies, 9, 85
- Ryle & Windram (1968) Ryle, M. & Windram, M. D. 1968, MNRAS, 138, 1
- Sales et al. (2007) Sales, L. V., Navarro, J. F., Abadi, M. G., & Steinmetz, M. 2007, MNRAS, 379, 1475
- Sanders et al. (2020) Sanders, J. S., Dennerl, K., Russell, H. R., et al. 2020, A&A, 633, A42
- Sato et al. (2005) Sato, K., Furusho, T., Yamasaki, N. Y., et al. 2005, PASJ, 57, 743
- Schwarz et al. (1992) Schwarz, R. A., Edge, A. C., Voges, W., et al. 1992, A&A, 256, L11
- Sheardown et al. (2019) Sheardown, A., Fish, T. M., Roediger, E., et al. 2019, ApJ, 874, 112
- Sijbring & de Bruyn (1998) Sijbring, D. & de Bruyn, A. G. 1998, A&A, 331, 901
- Simionescu et al. (2011) Simionescu, A., Allen, S. W., Mantz, A., et al. 2011, Science, 331, 1576
- Simionescu et al. (2012) Simionescu, A., Werner, N., Urban, O., et al. 2012, ApJ, 757, 182
- Springel (2005) Springel, V. 2005, MNRAS, 364, 1105
- Sunyaev et al. (2021) Sunyaev, R., Arefiev, V., Babyshkin, V., et al. 2021, A&A, 656, A132
- Tamura et al. (2014) Tamura, T., Yamasaki, N. Y., Iizuka, R., et al. 2014, ApJ, 782, 38
- van Weeren et al. (2024) van Weeren, R. J., Timmerman, R., Vaidya, V., et al. 2024, A&A, 692, A12
- Vikhlinin et al. (2005) Vikhlinin, A., Markevitch, M., Murray, S. S., et al. 2005, ApJ, 628, 655
- Walker et al. (2022) Walker, S. A., Mirakhor, M. S., ZuHone, J., et al. 2022, ApJ, 929, 37
- Zeldovich & Sunyaev (1969) Zeldovich, Y. B. & Sunyaev, R. A. 1969, Ap&SS, 4, 301
- Zhang et al. (2020) Zhang, C., Churazov, E., Dolag, K., Forman, W. R., & Zhuravleva, I. 2020, MNRAS, 498, L130
- Zhang et al. (2019) Zhang, C., Churazov, E., Forman, W. R., & Lyskova, N. 2019, MNRAS, 488, 5259
- Zhu et al. (2021) Zhu, Z., Simionescu, A., Akamatsu, H., et al. 2021, A&A, 652, A147
- Zhuravleva et al. (2015) Zhuravleva, I., Churazov, E., Arévalo, P., et al. 2015, MNRAS, 450, 4184
- Zuhone & Roediger (2016) Zuhone, J. A. & Roediger, E. 2016, Journal of Plasma Physics, 82, 535820301
Appendix A Chandra and XMM datasets
The OBSIDs of publicly available Chandra and XMM-Newton data used in this study are listed below.
For Chandra:
11713, 11714, 11715, 11716, 12025, 12033, 12036, 12037, 13989, 13990, 13991, 13992, 3209, 4289, 4946, 4947, 4948, 4949, 4950, 4951, 4952, 4953, 6139, 6145, 6146.
The initial data processing was done following the procedures described in Vikhlinin et al. (2005).
For XMM-Newton: 0085110101, 0085590201, 0151560101, 0204720101, 0204720201, 0305690101, 0305690301, 0305690401, 0305720101, 0305720301, 0305780101, 0405410101, 0405410201, 0673020201, 0673020301, 0673020401. The XMM-Newton data were cleaned for flares and then used similarly to Churazov et al. (2003).
Appendix B Radial X-ray surface brightness profile
As described in Sect. 2.3, eROSITA, Chandra, and XMM-Newton images were rebinned to the same spatial grid (2” pixels) and normalized to ensure a common flux calibration. A radial profile of the resulting image is shown in Fig. 12. The observed profile was approximated with a simple combination of a -model (the blue line in Fig. 12) and a constant (the dashed line). The very core (”) was excluded when fitting the profile. This model (sum of the -model and the constant) is used to remove a symmetric component from the X-ray image and express the residual deviations relative to the same model.